OpenStudy (anonymous):

solve 9^x = 3^(x - 2) for x

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

I'd be delighted to help you, but need to finish up helping 2 other students first.

OpenStudy (anonymous):

Would x = -2 ?

OpenStudy (mathmale):

You substitute x=-2 into the original equation abnd see wehther or not that equation is true. if it is, your result, x=-2, is correct.

OpenStudy (anonymous):

Okay, I substituted it and x = -2!

OpenStudy (anonymous):

2^x = 4^(x + 1) I think that one would be x = -2 also

OpenStudy (mathmale):

Actually, you replace every "x" with the value -2 and then deetermine whether or not the resulting equation is true. Try again, please. Or maybe I misread your response. Unsure.

OpenStudy (mathmale):

You got x=-2 already. All I'm asking you to do is to check that by subst. it back into the original equation.

OpenStudy (anonymous):

This one is a bit tricky for me 4^x = 10

OpenStudy (mathmale):

OK. Remember, I told you I'm completing work with 2 others so please be patient...I promise to be with you as soon as possible.

OpenStudy (mathmale):

solve 9^x = 3^(x - 2) for x What I see here is the following: x x-2 9 = 3 Hope that's reasonably clear for you. We're supposed to find the value of x, right?

OpenStudy (anonymous):

Well, I have found that one. X = -2 but this one is the one that I am stumped on 4^x = 10

OpenStudy (mathmale):

It'd be so much easier if these 2 different bases, 9 and 3, could be modified so that both are powers of the same base. Note that 2^2 = 4. Thus, 4 = 2^2. Basic examples, but perhaps these will help you to understand what I'm aiming at. take that 9 and rewrite it as 3^2. Substitute (3^2) for that 9 but copy everything else down as -is.

OpenStudy (mathmale):

Oh, you're finished with the previous problem and starting a new one? Sorry, I misunderstood.

OpenStudy (anonymous):

Yes, I am now trying to find 4^x = 10, solve for x

OpenStudy (mathmale):

I see you want to solve 4^x = 10 for x. That 4^x is an exponential function with base 4, isn't it?

OpenStudy (mathmale):

Can't remember, but you and I have probably talked about how the exponential function and the logarithmic function are inverses of each other. Do you recall that?

OpenStudy (anonymous):

I used a pogram to help me solve this, nd this is the answer I got: log(2) + log(5) ------------ 2log(2) I plugged it in and it was true, but I have to find a decimal and round to the nearest thousandths.

OpenStudy (mathmale):

My point, Ricky, is that if you take the (log to the base 4) of both sides of 4^x = 10, you'll end up with x = (log to the base 4) of 10.

OpenStudy (anonymous):

So would the answer I need be x = 1.6610 ?

OpenStudy (mathmale):

Please explain how you got that. Did you take the log to the base 4 of 10? If you can assure me that you did, then your answer is right. you should check it.

OpenStudy (anonymous):

Yes, and I checked it and it all checks out!

OpenStudy (mathmale):

Cool! Congratulations! It's great that you know the change of base formula. Ricky, I'd be glad to continue, but it's taken longer than I expected to finish helping the other student. Please start on a new problem; post it; do your best to get started and then tell me what y ou ne3ed to know next.

OpenStudy (mathmale):

Chosen a new problem yet?